OpenStudy (anonymous):

How long will it take 14 machines to complete a task, given that 6 machines can complete the same task in 7 days?

jagr2713 (jagr2713):

\[\huge \color {Blue}{Welcome \to OpenStudy!}\] @huntermolero

OpenStudy (anonymous):

\(\sf \Large \frac{14\ machines}{ x\ days}=\frac{6\ machines}{7\ days}\\x\ days=\Large \frac{14 (7)}{6}\)

OpenStudy (anonymous):

evaluate :)